# Improving Medium Range Severe Weather Prediction through Transformer Post-processing of AI Weather Forecasts

### v2

@@ 2.3.2 Verification Metrics
Brier $\text{Brier score}=\frac{1}{N}\sum_{t=1}^{N}(f_{t}-o_{t})^{2}$ (1)
@@
Brier skill $\text{Brier skill score}=1-\frac{BS_{f}}{BS_{ref}}$ (2)
where $BS_{f}$ is the Brier Score of the forecast model and $BS_{ref}$ is the Brier Score of a reference forecast. Here, we use the observation base rate as the reference forecast. Positive BSS values indicate improvement over the reference, with 1 representing perfect skill, 0 indicating no skill difference from the reference, and negative values indicating worse performance than the reference.
@@ 2.3.3 Explainable AI with Integrated Gradients
To gain insights into the decision-making process of our decoder-only transformer framework, we employ Explainable Artificial Intelligence (XAI) techniques, specifically Integrated Gradients (IG; Sundararajan et al. (2017)). IG is a feature attribution method that aims to explain the relationship between input features and model predictions by accumulating gradients along a linear path from a baseline input (the matrix of zeros represents the forecast mean state) to the actual input. In our application, we customize IG to diagnose feature sensitivity for specific forecast lead times, focusing on Day 3, 4, 5, and 6. For each of these days and grid points, we first filter test samples based on whether the models’ predicted severe weather probability from DOT(GFS+Convect), DOT(GFS), DOT(Pangu-GFS) and DOT(Pangu-HRES) all exceed a predefined threshold. The predefined threshold is based on the calibration curve shown in Fig. 5 with thresholds exceeding 0.3, 0.25, 0.2, and 0.15 for forecast Days 3, 4, 5, and 6 respectively. Then, for the samples that surpass this threshold for a given day, we compute the IG to attribute the prediction back to the input features. By aggregating these attributions across multiple samples for each target day, we can identify the key input forecast features that are most influential in the model’s prediction of severe weather events at different forecast horizons, thus providing interpretability to the transformer’s complex decision process.
@@ 3 Forecast Verification Period Overview
The forecast verification period used in this work (1 February 2024 – 22 July 2024) was a particularly active period for severe convective weather events within the CONUS. By the end of July 2024, the total number of tornado and wind reports in 2024 received at the Storm Prediction Center was in the top 3 among the last 15 severe weather seasons, while the number of hail reports was slightly below normal (SPC, 2025). Severe weather events occurring during April and May 2024 were particularly frequent and intense. During this two month period, the yearly total of tornado reports went from slightly below average in early April to near the 90th percentile of report totals over the 25 year period between 1999–2023 by the end of May (SPC, 2025). This two-month period included 11 days with $\geq$ 25 preliminary tornado reports across the CONUS and at least a slight risk of severe weather was issued every day somewhere over CONUS by the SPC, with 16/31 days having an enhanced risk, 5 days having a moderate risk, and 1 having a high risk. During the verification period, severe weather events were most common across the central Great Plains into the Midwest (i.e., eastern Colorado through Missouri and Indiana), with $\approx$10of days having $\geq$ 1 severe weather report across eastern Kansas and western Missouri (Fig. 2).
@@ 4.1 Forecast Skill versus Lead Time
Figure 3: Performance of various severe weather forecasting configurations as a function of forecast day on the test dataset, evaluated by (a) BSS, (b) PRAUC, and (c) ROCAUC. The lines in the bottom right legend represent different models or configuration. Specifically, GFS post-processed with a dense neural network is denoted as DNN, and the rest are decoder-only transformer (DOT) models. Statistical significance at the 95% confidence level was assessed using bootstrap resampling of 1,000 random sets of skill differences between model pairs at each lead time. Statistical significance was determined if the bootstrapped 95% confidence interval lies entirely above zero.
@@
Fig. 3 illustrates the performance of various severe weather forecasting configurations as a function of forecast day, evaluated using BSS (Fig. 3a), PRAUC (Fig. 3b), and ROCAUC (Fig. 3c). Notably, the GFS and GFS+Convect models trained with the DOT framework (DOT(GFS) and DOT(GFS+Convect)) consistently outperform the GFS models post-processed with DNN (DNN(GFS) and DNN(GFS+Convect)) across all forecast lead times and metrics. There is a significant performance drop for the DNN model when convective parameters are removed, as seen when comparing DNN(GFS) to DNN(GFS+Convect). Furthermore, comparing DOT(GFS) to both DNN(GFS) and DNN(GFS+Convect) shows that DOT(GFS) nearly doubles the BSS performance relative to DNN(GFS) for lead times less than 4 days, and also significantly outperforms DNN(GFS+Convect) after Day 1. This statistically significant improvement highlights the effectiveness of the transformer architecture in capturing temporal dependencies within the forecast sequence.
Additionally, DOT(Pangu-GFS) significantly outperforms DOT(GFS) in almost all lead times, and especially at Days 3–6 for all three metrics. Lastly, DOT(Pangu-HRES) significantly outperforms DOT(GFS) in all metrics at nearly all lead times except Day 8 and also outperforms DOT(GFS+Convect), across all three skill metrics from Day 3 to Day 6. DOT(Pangu-HRES) achieves a performance similar to that of DOT(Pangu-ERA5) after Day 5 and surpasses it by Day 8.
@@
Finally, the operational model mean (Mean(DOT(Pangu-GFS), DOT(Pangu-HRES), DOT(GFS+Convect))), representing the ensemble average output probability of ‘DOT(GFS+Convect)’, ‘DOT(Pangu-GFS)’, and ‘DOT(Pangu-HRES)’, consistently attains the best skill across all metrics and forecast days. This highlights the power of ensemble forecasting with potentially well-calibrated models, where averaging can effectively mitigate individual model errors and yield overall superior and more robust predictions. These results highlight the potential of using AI NWP emulators for medium-range severe weather forecasting. Our analysis in the following sections will focus solely on the verification of Day 3 to Day 6 forecasts where the differences between the GFS and Pangu-Weather forecasts are largest and statistically significant.
@@ 4.2 Sensitivity of Probability Threshold versus Lead Time
Comparing the models, DOT(Pangu-ERA5) often shows strong performance, particularly at shorter lead times (Day 3, Day 4) and higher thresholds Fig. 4c,d, maintaining relatively high precision and recall. DOT(Pangu-HRES) also demonstrates competitive performance, often closely following or slightly underperforming DOT(Pangu-ERA5) at shorter lead times but sometimes showing better retention of skill at Day 5 or Day 6 depending on the threshold. DOT(Pangu-GFS) generally exhibits lower performance than the Pangu-ERA5 and Pangu-HRES variants, especially at longer lead times. The baseline DOT(GFS) consistently shows the lowest performance among the models across most thresholds and lead times.
@@ 4.3 Reliability of Forecast Probability
As the forecast lead time increases from Day 3 to Day 6, two trends are apparent. First, the maximum forecast probability assigned by the models decreases. This is because as the lead time increases, the model is less confident with the higher probability, and also the models becomes less confident in their predictions, reflecting the increasing uncertainty associated with longer-range forecasts. Second, the reliability curves for all individual models tend to exhibit slight over-forecasting at Day 3, particularly for observed relative frequencies greater than 30%, but this diminishes somewhat at longer lead times. Interestingly, the DOT(GFS) model tends to struggle with predicting probability greater than 50% compared to other Pangu models.
@@
The Mean(DOT(Pangu-GFS), DOT(Pangu-HRES), DOT(GFS+Convect)) shows a different behavior. While it generally performs the best (as indicated by its superior Brier Skill Score in Fig. 3), it exhibits a tendency to under-forecast at higher probability intervals, especially for higher forecast probabilities. This under-confidence, while not ideal for perfect calibration, can be beneficial in terms of BSS because it avoids overly confident, incorrect predictions, which are heavily penalized by the Brier Score. In summary, while exhibiting some deviations from perfect calibration, all models produce generally reliable forecasts, and the operational model mean still produces the best BSS.
4.4 Seasonal Variation of Forecast Skill
@@
Figure 6: Seasonal evolution of the Brier Skill Score (BSS) difference relative to GFS for severe weather probability forecasts at lead times of (a) Day 3, (b) Day 4, (c) Day 5, and (d) Day 6. The BSS difference is calculated using a 6-week centered moving window applied to the samples, reducing sensitivity to daily forecast fluctuations. The x-axis represents the centered initialization time of the forecasts within the moving window. The y-axis represents the BSS difference, where positive values indicate improved skill compared to DOT(GFS).
@@
Here we investigate the seasonal evolution of the BSS difference relative to the DOT(GFS) model for Day 3 – 6 (Fig. 6a-d, respectively), calculated using a 6-week centered moving window. Several interesting patterns emerge from this time series analysis. First, across all lead times, most models demonstrate a positive BSS difference relative to DOT(GFS) for the majority of days, indicating improved skill that is persistent through the evaluation period. However, the magnitude of this improvement varies depending on models and lead time. For instance, DOT(Pangu-GFS) performs consistently worse than DOT(GFS) for a stretch of days in late March and early April 2024 at Day 6, and for shorter stretches at earlier lead-times. A case during this period will be examined in a later section. Secondly, there is a noticeable seasonal variation in the BSS difference. For instance, the BSS difference for many models tends to peak in late March and early April before gradually decreasing into the summer months. This fluctuation could be attributed to stronger synoptic forcing and more organized weather systems during the spring transition period, which may be better captured by the Pangu-Weather compared to GFS. Furthermore, the Pangu models exhibit considerable fluctuation in their BSS difference throughout the period, suggesting a sensitivity to specific forecast initialization times and potentially varying performance across different weather regimes. Periods of higher BSS difference might correspond to more active storm seasons or periods with stronger, more predictable synoptic-scale forcing. Conversely, periods of lower BSS difference or even negative differences could indicate challenges in specific weather situations or regimes where the strengths of the Pangu-Weather are less pronounced, or even where DOT(GFS) performs well. The Mean(DOT(Pangu-GFS), DOT(Pangu-HRES), DOT(GFS+Convect)), while also exhibiting temporal variability, generally maintains a more stable and consistently positive BSS difference, suggesting a degree of robustness achieved through ensemble averaging that mitigates some of the fluctuations seen in individual models.
@@ 4.5 Forecast Skill’s Dependence on Forecasted Convective Parameters
Figure 7: Brier Skill Score (BSS) and relative BSS difference conditioned on GFS-forecasted convective parameters. Panels (a-d) show the BSS of the DOT(GFS+Convect) model for Day 3 to 6, respectively, stratified by Surface-Based CAPE (SBCAPE) and surface to 500mb wind shear (SHEAR500) ranges. Panels (e-h) show the BSS difference between DOT(Pangu-GFS) and DOT(GFS+Convect) (DOT(Pangu-GFS) - DOT(GFS+Convect)) for the same days and convective parameter ranges. Panels (i-l) show the BSS difference between DOT(Pangu-HRES) and DOT(GFS+Convect) (DOT(Pangu-HRES) - DOT(GFS+Convect)). Positive values (red) indicate improved skill of Pangu models relative to GFS, while negative values (blue) indicate reduced skill. The number of forecast samples (N) within each SBCAPE/SHEAR500 bin is indicated in parentheses for the GFS model (panels a-d).
Figure 7 examines the BSS of severe weather forecasts conditioned on GFS-forecasted Surface-Based CAPE (SBCAPE) and surface to 500mb wind speed difference (SHEAR500) averaged over each lead time interval for Day 3 to 6. This allows us to see if the differences between the operational and AI models are a function of environmental characteristics. The first row (Fig. 7a-d) displays the BSS of the DOT(GFS+Convect) model across different ranges of SBCAPE and SHEAR500. Generally, for DOT(GFS+Convect), the BSS tends to increase with both increasing SBCAPE and SHEAR500, especially moving towards the highest SBCAPE range and higher shear categories. This indicates that GFS exhibits better skill in predicting severe weather when it forecasts stronger convective instability and wind shear, as expected in environments conducive to severe storms.
@@ 4.6 Feature Sensitivity
Figure 8: Integrated Gradients (IG) feature attribution heatmaps for Day 5 severe weather forecasts. Panels (a)-(d) show results for DOT(GFS+Convect), and panels (e)-(g) show results for DOT(Pangu-GFS). Each heatmap displays the attribution of input features (y-axis) across the 8-day input sequence (x-axis). Features within each group (Upper, Surface, Static, Convective) are sorted vertically by the absolute magnitude of their Day 5 attribution, from highest to lowest importance. The overall importance rank for each feature within its model is shown in parentheses. Red cells indicate that an increase in the feature value contributes to a higher predicted probability, while blue indicates a contribution to a lower probability. Attributions are averaged on common cases in the test period with relatively with probability threshold exceeding 0.2 on Day 5 for DOT(GFS+Convect), DOT(GFS), DOT(Pangu-GFS) and DOT(Pangu-HRES). See supplementary material Fig. S1 for other models and days.
Figure 8 presents an example of Integrated Gradients (IG) attribution heatmaps comparing DOT(GFS+Convect) (Fig. 8a-d) and DOT(Pangu-GFS) (Fig. 8e-g) for Day 5 severe weather forecasts, with additional model and day comparisons available in the supplementary material (Fig. S1). Consistent with the transformer architecture using a causal mask, the features corresponding to the target forecast day (Day 5, the 5th column in each heatmap) generally exhibit the largest attribution magnitudes for both models. This signifies the strong reliance on the atmospheric conditions forecast for the specific target day. However, features from preceding days (Days 1-4) also show non-zero attributions, indicating that the model leverages the temporal evolution of the environment leading up to Day 5. Attributions for subsequent days (Days 6-8) are minimal, reflecting the causal mask preventing the use of future information. Furthermore, static variable like the day of lead time for forecasting is ranked in the top 5 important features in both models.
@@
In contrast, DOT(GFS+Convect) heavily relies on surface temperature (t2m, rank 1) and explicitly uses convective parameters like surface-based CAPE (sbcape, rank 3) which shows a strong positive attribution on Day 5. Interestingly, 2m dewpoint (q2m, rank 6) shows a counterintuitive pattern: positive attribution in the days leading up to Day 5, switching to negative attribution on Day 5 itself, potentially reflecting model behavior related to potential ongoing convection. While convective parameters like SBCAPE and precipitation are, unsurprisingly, important in GFS+Convect, 0-3km storm relative helicity (SRH03) exhibits less influence than expected based on prior studies (e.g. Kerr and Darkow (1996)) linking it to supercell risk. A consistent, yet counterintuitive, pattern emerges in models for features such as q2m, z200, and u850 where the positive attribution preceding the forecast day shifts to the negative attribution on the forecast day. This attribution behavior is consistent across most of the forecast models and lead-times. For u850, stronger negative zonal winds may indicate more strongly backed flow or better moisture advection on the day of a severe weather event, but in the days prior, a more westerly zonal wind may be more conducive for severe weather during the following days. Other fields, such as q2m in the GFS model have a similar behavior, and may reflect systematic variations in moisture due to the influence of ongoing convection in the model during the forecast day.

### v3

@@ 2.3.2 Verification Metrics
$\text{Brier score}=\frac{1}{N}\sum_{t=1}^{N}(f_{t}-o_{t})^{2}$ (1)
@@
$\text{Brier skill score}=1-\frac{BS_{f}}{BS_{ref}}$ (2)
where $BS_{f}$ is the Brier Score of the forecast model and $BS_{ref}$ is the Brier Score of a reference forecast. Here, the reference forecast $BS_{ref}$ is the Brier Score of a simple climatological forecast, where the forecast probability is the static, overall frequency of severe weather events (the base rate) in the provided test period dataset or period. Positive BSS values indicate improvement over the reference, with 1 representing perfect skill, 0 indicating no skill difference from the reference, and negative values indicating worse performance than the reference.
@@ 2.3.3 Explainable AI with Integrated Gradients
To gain insights into the decision-making process of our decoder-only transformer framework, we employ Explainable Artificial Intelligence (XAI) techniques, specifically Integrated Gradients (IG; Sundararajan et al. (2017)). IG is a feature attribution method that aims to explain the relationship between input features and model predictions by accumulating gradients along a linear path from a baseline input (the matrix of zeros represents the forecast mean state) to the actual input. In our application, we customize IG to diagnose feature sensitivity for specific forecast lead times, focusing on Day 3, 4, 5, and 6. For each of these days and grid points, we first filter test samples based on whether the models’ predicted severe weather probability from DOT(GFS+Convect), DOT(GFS), DOT(Pangu-GFS) and DOT(Pangu-HRES) all exceed a predefined threshold. The predefined threshold is based on the calibration curve shown in Fig. 6 with thresholds exceeding 0.3, 0.25, 0.2, and 0.15 for forecast Days 3, 4, 5, and 6 respectively. Then, for the samples that surpass this threshold for a given day, we compute the IG to attribute the prediction back to the input features. The attributions are aggregated across the 6-hourly time steps within each 24-hour forecast period, providing daily-scale feature importance rather than subdaily temporal resolution. By aggregating these attributions across multiple samples for each target day, we can identify the key input forecast features that are most influential in the model’s prediction of severe weather events at different forecast horizons, thus providing interpretability to the transformer’s complex decision process.
@@ 3 Forecast Verification Period Overview
The forecast verification period used in this work (1 February 2024 – 22 July 2024) was a particularly active period for severe convective weather events within the CONUS. By the end of July 2024, the total number of tornado and wind reports in 2024 received at the Storm Prediction Center was in the top 3 among the last 15 severe weather seasons, while the number of hail reports was slightly below normal (SPC, 2025). Severe weather events occurring during April and May 2024 were particularly frequent and intense. During this two month period, the yearly total of tornado reports went from slightly below average in early April to near the 90th percentile of report totals over the 25 year period between 1999–2023 by the end of May (SPC, 2025). This two-month period included 11 days with $\geq$ 25 preliminary tornado reports across the CONUS and at least a slight risk of severe weather was issued every day somewhere over CONUS by the SPC, with 16/31 days having an enhanced risk, 5 days having a moderate risk, and 1 having a high risk. During the verification period, severe weather events were most common across the central Great Plains into the Midwest (i.e., eastern Colorado through Missouri and Indiana), with $\approx$10% of days having $\geq$ 1 severe weather report across eastern Kansas and western Missouri (Fig. 2).
@@ 4.1 Forecast Skill versus Lead Time
4.1.1 Transformer (DOT) versus DNN Architecture
@@
A direct comparison between the Decoder-only Transformer (DOT) and the Dense Neural Network (DNN) architectures, both utilizing GFS inputs, reveals the significant advantages of the transformer-based approach (Fig. 3). The DOT(GFS+Convect) and DOT(GFS) models consistently outperform their DNN counterparts (DNN(GFS+Convect) and DNN(GFS)) across all forecast lead times and performance metrics. Notably, there is a significant performance drop for the DNN model when convective parameters are removed, as seen when comparing DNN(GFS) to DNN(GFS+Convect). In contrast, the DOT(GFS) model, even without explicit convective parameters, significantly outperforms both DNN models. For lead times less than 4 days, DOT(GFS) nearly doubles the BSS performance relative to DNN(GFS) and also significantly outperforms DNN(GFS+Convect) after Day 1. This statistically significant improvement highlights the effectiveness of the transformer architecture in capturing temporal dependencies within the forecast sequence.
Figure 3: The plots show (a) Base Rate BSS, (b) PRAUC, and (c) ROCAUC performance comparison as a function of forecast day between DOT and DNN architectures using GFS forecast. The lines in the bottom right legend represent different models or configuration. Statistical significance at the 95% confidence level was assessed for the specific model pairs listed in the legend below the plot. For each pair, 1,000 bootstrap resamples of the skill difference were computed at each lead time. A symbol is shown if the corresponding bootstrapped 95% confidence interval lies entirely above zero, indicating a statistically significant improvement.
4.1.2 AI NWP (Pangu-Weather) versus NWP (GFS) Forecast
When comparing the performance of the DOT architecture with different forecast sources—AI-driven Pangu-Weather forecasts versus traditional NWP GFS forecasts—the results demonstrate the potential capability of the AI-based forecast, particularly in the medium-range (Fig. 4). The DOT(Pangu-GFS) model significantly outperforms the DOT(GFS) model in almost all lead times, with the most substantial improvements observed between Days 3 and 6 across all three metrics. Lastly, DOT(Pangu-HRES) significantly outperforms DOT(GFS) in all metrics at nearly all lead times except Day 8 and also outperforms DOT(GFS+Convect), across all three skill metrics from Day 3 to Day 6. DOT(Pangu-HRES) achieves a performance similar to that of DOT(Pangu-ERA5) after Day 5 and surpasses it by Day 8. These findings underscore the value of using AI NWP emulators like Pangu-Weather for generating more skillful medium-range severe weather forecasts.
Figure 4: Similar to Fig. 3, this figure compares the performance of the DOT post-processing architecture, but using Pangu-Weather forecasts versus GFS forecasts as input.
@@ 4.1.2 AI NWP (Pangu-Weather) versus NWP (GFS) Forecast
Finally, the operational model mean (Mean(DOT(Pangu-GFS), DOT(Pangu-HRES), DOT(GFS+Convect))), representing the ensemble average output probability of ‘DOT(GFS+Convect)’, ‘DOT(Pangu-GFS)’, and ‘DOT(Pangu-HRES)’, consistently attains the best skill across all metrics and forecast days. This highlights the power of ensemble forecasting with potentially well-calibrated models, where averaging can effectively mitigate individual model errors and yield overall superior and more robust predictions. Our analysis in the following sections will focus solely on the verification of Day 3 to Day 6 forecasts where the differences between the GFS and Pangu-Weather forecasts are largest and statistically significant.
@@ 4.2 Sensitivity of Probability Threshold versus Lead Time
Comparing the models, DOT(Pangu-ERA5) often shows strong performance, particularly at shorter lead times (Day 3, Day 4) and higher thresholds (Fig. 5c,d), maintaining relatively high precision and recall. DOT(Pangu-HRES) also demonstrates competitive performance, often closely following or slightly underperforming DOT(Pangu-ERA5) at shorter lead times but sometimes showing better retention of skill at Day 5 or Day 6 depending on the threshold. DOT(Pangu-GFS) generally exhibits lower performance than the Pangu-ERA5 and Pangu-HRES variants, especially at longer lead times. The baseline DOT(GFS) consistently shows the lowest performance among the models across most thresholds and lead times.
@@ 4.3 Reliability of Forecast Probability
As the forecast lead time increases from Day 3 to Day 6, two trends are apparent. First, the maximum forecast probability assigned by the models decreases. This is because as the lead time increases, the model is less confident with the higher probability, and also the models become less confident in their predictions, reflecting the increasing uncertainty associated with longer-range forecasts. Second, the reliability curves for all individual models tend to exhibit slight over-forecasting at Day 3, particularly for observed relative frequencies greater than 30%, but this diminishes somewhat at longer lead times. Interestingly, the DOT(GFS) model tends to struggle with predicting probability greater than 50% compared to other Pangu models.
@@
The Mean(DOT(Pangu-GFS), DOT(Pangu-HRES), DOT(GFS+Convect)) shows a different behavior. While it generally performs the best (as indicated by its superior Brier Skill Score in Fig. 4), it exhibits a tendency to under-forecast at higher probability intervals. This under-confidence is not ideal for perfect calibration, it can be beneficial for the overall Brier Skill Score (BSS). Due to the Brier Score’s quadratic nature, it penalizes large errors disproportionately. For example, an incorrect forecast of 90% receives more than double the penalty of an incorrect 60% forecast. By systematically avoiding high-confidence predictions, the model mitigates the risk of these large, heavily penalized errors, which often improves the average score. In summary, while exhibiting some deviations from perfect calibration, all models produce generally reliable forecasts, and the operational model mean still produces the best BSS.
4.4 Forecast Skill Time Series
@@
Figure 7: Time series of the Brier Skill Score (BSS) difference relative to DOT(GFS) for severe weather probability forecasts at lead times of (a) Day 3, (b) Day 4, (c) Day 5, and (d) Day 6. The BSS difference is calculated using a 6-week centered moving window applied to the samples, reducing sensitivity to daily forecast fluctuations. The x-axis represents the centered initialization time of the forecasts within the moving window. The y-axis represents the BSS difference, where positive values indicate improved skill compared to DOT(GFS).
@@
Here we investigate the temporal evolution of the BSS difference relative to the DOT(GFS) model for Day 3 – 6 (Fig. 7a-d, respectively), calculated using a 6-week centered moving window. Several interesting patterns emerge from this time series analysis. First, across all lead times, most models demonstrate a positive BSS difference relative to DOT(GFS) for the majority of days, indicating improved skill that is persistent through the evaluation period. However, the magnitude of this improvement varies depending on models and lead time. For instance, DOT(Pangu-GFS) performs consistently worse than DOT(GFS) for a stretch of days in late March and early April 2024 at Day 6, and for shorter stretches at earlier lead-times. A case during this period will be examined in a later section. Secondly, there is a noticeable temporal variation in the BSS difference. For instance, the BSS difference for many models tends to peak in late March and early April before gradually decreasing into the summer months. This fluctuation could be attributed to stronger synoptic forcing and more organized weather systems during the spring transition period, which may be better captured by the Pangu-Weather compared to GFS. Furthermore, the Pangu models exhibit considerable fluctuation in their BSS difference throughout the period, suggesting a sensitivity to specific forecast initialization times and potentially varying performance across different weather regimes. Periods of higher BSS difference might correspond to more active storm seasons or periods with stronger, more predictable synoptic-scale forcing. Conversely, periods of lower BSS difference or even negative differences could indicate challenges in specific weather situations or regimes where the strengths of the Pangu-Weather are less pronounced, or even where DOT(GFS) performs well. The Mean(DOT(Pangu-GFS), DOT(Pangu-HRES), DOT(GFS+Convect)), while also exhibiting temporal variability, generally maintains a more stable and consistently positive BSS difference, suggesting a degree of robustness achieved through ensemble averaging that mitigates some of the fluctuations seen in individual models. Similar findings are also supported by the data in Fig. S2, where DOT(GFS+Convect) is used as the baseline.
@@ 4.5 Forecast Skill’s Dependence on Forecasted Convective Parameters
Figure 8: Brier Skill Score (BSS) and relative BSS difference conditioned on GFS-forecasted convective parameters. Panels (a-d) show the BSS of the DOT(GFS+Convect) model for Day 3 to 6, respectively, stratified by Surface-Based CAPE (SBCAPE) and surface to 500mb wind shear (SHEAR500) ranges. Panels (e-h) show the BSS difference between DOT(Pangu-GFS) and DOT(GFS+Convect) (DOT(Pangu-GFS) - DOT(GFS+Convect)) for the same days and convective parameter ranges. Panels (i-l) show the BSS difference between DOT(Pangu-HRES) and DOT(GFS+Convect) (DOT(Pangu-HRES) - DOT(GFS+Convect)). Positive values (red) indicate improved skill of Pangu models relative to GFS, while negative values (blue) indicate reduced skill. The number of forecast samples (N) within each SBCAPE/SHEAR500 bin is indicated in parentheses for the GFS model (panels a-d). Pair-wise significance testing similar to Fig. 3 to compare against DOT(Pangu-GFS) and DOT(Pangu-HRES) with DOT(GFS+Convect) for different forecasted convective parameters regimes. With asterisks mark after score indicates statistically significantly better than DOT(GFS+Convect).
Figure 8 examines the BSS of severe weather forecasts conditioned on GFS-forecasted SBCAPE and surface to 500mb wind speed difference (SHEAR500) averaged over each lead time interval for Day 3 to 6. This allows us to see if the differences between the operational and AI models are a function of environmental characteristics. The first row (Fig. 8a-d) displays the BSS of the DOT(GFS+Convect) model across different ranges of SBCAPE and SHEAR500. Generally, for DOT(GFS+Convect), the BSS tends to increase with both increasing SBCAPE and SHEAR500, especially moving towards the highest SBCAPE range and higher shear categories. This indicates that GFS exhibits better skill in strongly-forced synoptic regimes, where its performance is less reliant on the intricacies of its convective parameterization scheme.
@@ 4.6 Feature Sensitivity
Figure 9: Integrated Gradients (IG) feature attribution heatmaps for Day 5 severe weather forecasts. Panels (a)-(d) show results for DOT(GFS+Convect), and panels (e)-(g) show results for DOT(Pangu-GFS). Each heatmap displays the attribution of input features (y-axis) across the 8-day input sequence (x-axis). Features within each group (Upper, Surface, Static, Convective) are sorted vertically by the absolute magnitude of their Day 5 attribution, from highest to lowest importance. The overall importance rank for each feature within its model is shown in parentheses. Red cells indicate that an increase in the feature value contributes to a higher predicted probability, while blue indicates a contribution to a lower probability. Attributions for future predictors (Days 6–8) are effectively zero due to the causal mask, with the small, non-zero IG values arising from numerical artifacts in masked-attention softmax precision and IG integration. Attributions are averaged on common cases in the test period with relatively with probability threshold exceeding 0.2 on Day 5 for DOT(GFS+Convect), DOT(GFS), DOT(Pangu-GFS) and DOT(Pangu-HRES). See supplementary material Fig. S3 for other models and days.
Figure S3 presents an example of Integrated Gradients (IG) attribution heatmaps comparing DOT(GFS+Convect) (Fig. S3a-d) and DOT(Pangu-GFS) (Fig. S3e-g) for Day 5 severe weather forecasts, with additional model and day comparisons available in the supplementary material (Fig. S3). Consistent with the transformer architecture using a causal mask, the features corresponding to the target forecast day (Day 5, the 5th column in each heatmap) generally exhibit the largest attribution magnitudes for both models. This signifies the strong reliance on the atmospheric conditions forecast for the specific target day. However, features from preceding days (Days 1-4) also show non-zero attributions, indicating that the model leverages the temporal evolution of the environment leading up to Day 5. Attributions for subsequent days (Days 6–8) are effectively zero, reflecting the causal mask that prevents the model from using future predictors. The small, non-zero IG values are due to numerical noise from masked-attention softmax precision limits, floating-point accumulation in IG integration, and the fact that inference is performed with all Days 1–8 predictors present. These artifacts are many orders of magnitude smaller than the relevant attributions and have no impact on interpretation. Lastly, static variable like the day of lead time for forecasting is ranked in the top 5 important features in both models.
@@
In comparison, while DOT(GFS+Convect) also heavily relies on surface temperature (t2m, rank 1), it can additionally and explicitly use convective parameters like surface-based CAPE (sbcape, rank 3) which shows a strong positive attribution on Day 5. Interestingly, 2m specific humidity (q2m, rank 6) shows a counterintuitive pattern: positive attribution in the days leading up to Day 5, switching to negative attribution on Day 5 itself, potentially reflecting model behavior related to potential ongoing convection. While convective parameters like SBCAPE and precipitation are, unsurprisingly, important in GFS+Convect, 0-3km storm relative helicity (SRH03) exhibits less influence than expected based on prior studies (e.g. Kerr and Darkow (1996)) linking it to supercell risk. A consistent, yet counterintuitive, pattern emerges in models for features such as q2m, z200, and u850 where the positive attribution preceding the forecast day shifts to the negative attribution on the forecast day. This attribution behavior is consistent across most of the forecast models and lead-times. For u850, stronger negative zonal winds may indicate more strongly backed flow or better moisture advection on the day of a severe weather event, but in the days prior, a more westerly zonal wind may be more conducive for severe weather during the following days. Other fields, such as q2m in the GFS model have a similar behavior, and may reflect systematic variations in moisture due to the influence of ongoing convection in the model during the forecast day.